For the most recent year available, the mean annual cost to attend a private

university in the United States was $26,889. Assume the distribution of annual costs follows the

normal probability distribution and the standard deviation is $4,500. Ninety-five percent of all

students at private universities pay less than what amount?


1
Expert's answer
2021-07-13T09:49:35-0400

Let "X=" annual cost: "X\sim N(\mu, \sigma^2)."

Given "\mu=26889, \sigma=4500."


"P(X<\dfrac{x-26889}{4500})=0.95"

"\dfrac{x-26889}{4500}=1.6449"

"x=4500(1.6449)+26889"

"x=\$34291"

Ninety-five percent of all students at private universities pay less than $34,291.